# Non-stationary Spatio-Temporal Modeling Using the Stochastic Advection-Diffusion Equation

### v1

@@ 3.4 Discrete hierarchical model
${\bm{w}}=\begin{bmatrix}{\bm{w}}^{0}\\ \vdots\\ {\bm{w}}^{T-1}\end{bmatrix}$
be the vector of the discrete solution of the stochastic advection-diffusion-reaction equation at all time grid points. For ease of description, we assume that all observation time points match one of the time grid points. I.e., $\{t_{1},\ldots,t_{n}\}\subset\{t_{0}^{\mathrm{G}},\ldots,t_{T-1}^{\mathrm{G}}\}$. In the simulation study and the application, this will always be the case.
We can then write the model in vector form,
$\bm{y}=\mathbf{X}\bm{\beta}+\mathbf{E}{\bm{w}}+\bm{\epsilon},$ (13)
where $\mathbf{X}$ is a design matrix of the covariates, where each row $i$ is $\bm{x}(\bm{s}_{i},t_{i})^{\mathrm{T}}$, $\bm{\beta}$ is the coefficients of the covariates, $\bm{\epsilon}\sim{\cal N}(\bm{0},\sigma^{2}_{n}\mathbf{I})$, and $\mathbf{E}$ is a $n\times KT$ matrix that selects the correct grid cell in the correct time grid point for each observation. Note that $\mathbf{E}$ has exactly one $1$ on each row and all other elements are 00.
@@
Stage 1: $\bm{y}\mid\bm{\beta},{\bm{w}},\sigma_{n}^{2}\sim{\cal N}\left(\mathbf{X}\bm{\beta}+\mathbf{E}{\bm{u}},\sigma_{n}^{2}\mathbf{I}\right)$
Stage 2: ${\bm{w}}\mid\bm{\theta}\sim{\cal N}(\bm{0},\mathbf{Q}^{-1})\enspace,\quad\bm{\beta}\sim{\cal N}\left(\bm{0},V_{\beta}\mathbf{I}\right),$
where details about $\mathbf{Q}$ is given in Appendix A.. Since ${\bm{w}}$ and $\beta$ are jointly Gaussian given $\bm{\theta}$, we collect them together as ${\bm{z}}=[{\bm{w}},\bm{\beta}]$, and also let $\mathbf{S}=[\mathbf{E},\mathbf{X}]$ such that the hierarchical model becomes
Stage 1: $\bm{y}\mid{\bm{z}},\sigma_{n}^{2}\sim{\cal N}\left(\mathbf{S}{\bm{z}},\sigma_{n}^{2}\mathbf{I}\right)$
@@
If the parameters $\bm{\theta}$ and $\sigma_{n}^{2}$ are known, then the conditional distribution of the latent field is
${\bm{z}}|\bm{y},\bm{\theta},\sigma_{n}^{2}\sim\mathcal{N}(\bm{\mu}_{\mathrm{C}},\mathbf{Q}_{\mathrm{C}}^{-1}),$ (14)
@@
$\mathbf{Q}_{\mathrm{C}}=\mathbf{Q}+\mathbf{S}^{\top}\mathbf{S}\sigma^{-2}_{n},$ (15)
@@
$\bm{\mu}_{\mathrm{C}}=\mathbf{Q}_{\mathrm{C}}^{-1}\mathbf{S}^{\top}\bm{y}\sigma^{-2}_{n}.$ (16)
Using this conditional distribution, the point prediction for an unobserved location $\bm{s}^{*}\in\mathcal{D}$ is given by
$\mathbb{E}\left[\bm{y}^{*}|\bm{y},\sigma_{n}^{2},\bm{\theta}\right]=\mathbf{S}_{\mathrm{P}}\bm{\mu}_{\mathrm{C}},$
where $\mathbf{S}_{\mathrm{P}}$ is the corresponding design matrix for the prediction location. The predictive distribution $\bm{y}^{*}|\bm{y},\sigma_{n}^{2},\bm{\theta}$ is Gaussian, and the variance can be found either through algorithms that compute a partial inverse of $\mathbf{Q}_{\mathrm{C}}$ (Rue and Held,, 2010) or estimated through simulation. When forecasts are desired, one includes additional time steps in the discretization that are unobserved, and use the same approach as above.
@@ 3.5 Parameter inference
$\log\pi(\bm{\theta},\sigma^{2}_{n}|\bm{y})$ $=\mathrm{Const}+\log\pi(\bm{\theta},\sigma^{2}_{n})+\frac{1}{2}\log|\mathbf{Q}_{z}|-\frac{1}{2}\log|\mathbf{Q}_{\mathrm{C}}|+$ (17)
$\phantom{=}-\frac{n}{2}\log\sigma^{2}_{n}-\frac{1}{2}\bm{\mu}_{\mathrm{C}}^{\mathrm{T}}\mathbf{Q}_{z}\bm{\mu}_{\mathrm{C}}-\frac{1}{2\sigma^{2}_{n}}(\bm{y}-\mathbf{S}\bm{\mu}_{\mathrm{C}})^{\top}(\bm{y}-\mathbf{S}\bm{\mu}_{\mathrm{C}}),$
where
and we use $\pi(\bm{\theta},\sigma_{n}^{2})\propto 1$ in this paper since we are emulating densely observed numerical model output and do not need extra regularization. For the derivation of the log-likelihood, we refer the reader to Berild and Fuglstad, (2023, Appendix A.4).
Inference is done using some optimization algorithm on the log-posterior to find the parameters that maximize it. To aid the optimization, the gradient of the log-posterior is often used. However, calculating the analytic gradient of the log-posterior requires calulcating $\mathrm{Tr}((\mathbf{Q}_{z}^{-1}-\mathbf{Q}_{\mathrm{C}}^{-1})\frac{\partial\mathbf{Q}_{z}}{\partial\theta_{i}})$ where $\mathrm{Tr}(\cdot)$ denotes the trace of the matrix. Calculation of the inverse of the precision matrices can be avoided by computing a partial inverse only for the non-zero elements (Rue and Held,, 2010), but it is still quite computationally expensive. An alternative, to speed up computations further, is to use the Hutchinson estimator (Hutchinson,, 1990) which is a stochastic approximation of the trace of the matrix,
$\mathrm{Tr}(\mathbf{A})=\mathbb{E}\left[\bm{\xi}^{\top}\mathbf{A}\bm{\xi}\right]\approx\frac{1}{n}\sum_{i=1}^{n}\bm{\xi}_{i}^{\top}\mathbf{A}\bm{\xi}_{i},$
where $\bm{\xi}_{i}$ are independent samples from the Rademacher distribution, i.e., $\bm{\xi}_{i}=\pm 1$ with equal probability. This is much faster since it only requires $n$ matrix-vector products and solves, but it makes the optimization problem stochastic.
The drawback of stochastic optimization is that the gradient is noisy, and this can lead to erratic behavior in the optimization, but it is often used because the computational speedup is so significant that the convergence is still faster. There is a vast literature on stochastic gradient methods, and we refer the reader to Ruder, (2017) for an overview. Many of these methods tackle the problem of noisy gradients and use techniques such as momentum, adaptive learning rates, and averaging of gradients to reduce the noise. In this work, we use the Adam optimizer (Kingma and Ba,, 2017) which is a popular choice for optimization of neural networks and other large models.
@@ 4.1 True process and observation model
For simplicity, the dampening coefficient is constant $\kappa(\cdot)\equiv\exp(-2)$, the isotropic diffusion parameter is $\gamma(\cdot)\equiv\exp(-1)$, and the scaling parameter for the forcing is $\tau=\exp(-4)$. This defines the true spatio-temporal process $\{u_{\mathrm{true}}(\bm{s},t):\bm{s}\in\mathcal{D},t\in\mathcal{T}\}$, which we call NStat-True. The boundaries of the spatial domain are set to have zero flow going in or out of the domain. Furthermore, in order to reduce the effect of the boundary conditions, a buffer of $5$ grid cells is added to each side of the domain $\mathcal{D}$ resulting in a new domain $\mathcal{D}_{\mathrm{B}}$ of $M=60$ and $N=60$ grid cells. We also apply an advection tapering within this buffer, where the advection vector field is scaled linearly to zero from the boundary of $\mathcal{D}$ to the boundary of $\mathcal{D}_{\mathrm{B}}$. This is to avoid strong advection that suddenly stops at the boundary and inflates marginal variances strongly.
@@
$y_{i}=u_{\mathrm{true}}(\bm{s}_{i},t_{i})+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{n}^{2})\quad\forall i.$ (20)
Here, the noise is i.i.d. Gaussian with variance $\sigma_{n}^{2}=0.001$. Figure 3 shows one realization, marginal variance, and spatio-temporal correlation with the marked location.
@@ 4.2 Goals, and training and test data
We aim to mimic the emulation of a numerical model that is densely observed with a high signal-to-noise ratio. From NStat-True, we observe all grid cells except the grid cells in the rectangle $[6.1,11.6]\times[3.7,9.5]\subset\mathcal{D}$ (rounded to the first decimal), to assess the ability to reconstruct the missing rectangle using data up to the given time, and to forecast for different lags forward in time. The observed grid cells are observed with i.i.d. Gaussian measurement noise with nugget variance $\sigma_{n}^{2}=0.001$.
We generate $20$ realization of the NStat-True for training data that excludes the rectangular area and $20$ realization for the test data that is containing the rectangular area. The training data is demonstrated in Figure 3a. We aim to assess accuracy in reconstructing the masked area from the surrounding observations, and accuracy in forecasting the entire field for different lags forward in time. Let a point prediction for location $\bm{s}_{i}$ be denoted $\hat{y}^{*}_{i,j}=\mathbb{E}[y^{*}_{i}|\bm{y}_{j},\hat{\sigma_{n}^{2}},\hat{\bm{\theta}}]$, where $\bm{y}_{j}$ is the observed data of the $j$th test set, and $\hat{\sigma_{n}^{2}}$ and $\hat{\bm{\theta}}$ are the estimated model parameters. Point predictions are evaluated using root mean square error (RMSE)
$\mathrm{RMSE}(\Delta k_{p},j)=\sqrt{\frac{1}{n}\sum_{i=1}^{n}\left(y_{i}^{*}-\hat{y}^{*}_{i,j}\right)^{2}},$
where $n$ is the number of prediction location and $\Delta k_{p}=k_{p}-k_{o}\in\{0,1,2,\ldots,11\}$ is the difference in between predicted $k_{p}$ and observed $k_{o}$ timesteps where $k_{o},k_{p}\in\{1,2,\ldots,12\}$. We also assess the predictive distributions using the continuous ranked probability score (CRPS, Gneiting and Raftery, (2007))
$\mathrm{CRPS}(\Delta k_{p},j)=$ $\frac{1}{n}\sum_{i=1}^{n}\sigma_{i,j}(z_{i,j}(2\Phi(z_{i,j})-1)+2\phi(z_{i,j})-\pi^{-1/2}),\quad\textrm{for}$
$z_{i,j}=$ $\frac{y_{i}^{*}-\hat{y}^{*}_{i,j}}{\sigma_{i,j}},$
where $\sigma_{i,j}$ is the standard deviation of the predictive distribution $\mathrm{Var}(y^{*}_{i}|\bm{y}_{-j},\hat{\sigma_{n}^{2}},\hat{\bm{\theta}})$, $\Phi(\cdot)$ is the cumulative distribution function of the standard normal distribution, and $\phi(\cdot)$ is the density function of the standard normal distribution. Statistics, such as the average and standard deviation, for these metrics are calculated for each $\Delta k_{p}$ and candidate model over all $j=1,2,\ldots,20$ realizations in the test data.

### v2

@@ 3.4 Discrete hierarchical model
${\bm{u}}=\begin{bmatrix}{\bm{u}}^{0}\\ \vdots\\ {\bm{u}}^{T-1}\end{bmatrix}$
be the vector of the discrete solution of the stochastic advection-diffusion-reaction equation at all time grid points. We assume a piece-wise constant approximation so that for an arbitrary location $\bm{s}\in\mathcal{D}$ and time point $t\in\mathcal{T}$, the location is allocated to its corresponding grid cell in the spatial discretization, and the time point is allocated to its corresponding half open time interval $t\in[t_{k}^{\mathrm{G}},t_{k+1}^{\mathrm{G}})$ for the appropriate $k$. For ease of description, we assume that all observation time points match one of the time grid points. I.e., $\{t_{1},\ldots,t_{n}\}\subset\{t_{0}^{\mathrm{G}},\ldots,t_{T-1}^{\mathrm{G}}\}$. In the simulation study and the application, this will always be the case.
The model can then be written in vector form,
$\bm{y}=\mathbf{X}\bm{\beta}+\mathbf{E}{\bm{u}}+\bm{\epsilon},$ (13)
where $\mathbf{X}$ is a design matrix of the covariates, where each row $i$ is $\bm{x}(\bm{s}_{i},t_{i})^{\mathrm{T}}$, $\bm{\beta}$ is the coefficients of the covariates, $\bm{\epsilon}\sim{\cal N}(\bm{0},\sigma^{2}_{\mathrm{N}}\mathbf{I})$, and $\mathbf{E}$ is a $n\times KT$ matrix that selects the correct grid cell in the correct time grid point for each observation. Note that $\mathbf{E}$ has exactly one $1$ on each row and all other elements are 00.
@@
Stage 1: $\bm{y}\mid\bm{\beta},{\bm{u}},\sigma_{\mathrm{N}}^{2}\sim{\cal N}\left(\mathbf{X}\bm{\beta}+\mathbf{E}{\bm{u}},\sigma_{\mathrm{N}}^{2}\mathbf{I}\right)$
Stage 2: ${\bm{u}}\mid\bm{\theta}\sim{\cal N}(\bm{0},\mathbf{Q}^{-1})\enspace,\quad\bm{\beta}\sim{\cal N}\left(\bm{0},V_{\beta}\mathbf{I}\right),$
where $\mathbf{Q}$ is a $KT\times KT$ block tridiagonal matrix with sparse blocks of size $K\times K$ as described in Appendix A. Since ${\bm{w}}$ and $\beta$ are jointly Gaussian given $\bm{\theta}$, we collect them together as ${\bm{z}}=[{\bm{u}}^{\mathrm{T}},\bm{\beta}^{\mathrm{T}}]^{\mathrm{T}}$, and also let $\mathbf{S}=[\mathbf{E},\mathbf{X}]$ such that the hierarchical model becomes
Stage 1: $\bm{y}\mid{\bm{z}},\sigma_{\mathrm{N}}^{2}\sim{\cal N}\left(\mathbf{S}{\bm{z}},\sigma_{\mathrm{N}}^{2}\mathbf{I}\right)$
@@
If the parameters $\bm{\theta}$ and $\sigma_{\mathrm{N}}^{2}$ are known, and
then the conditional distribution of the latent field is
${\bm{z}}|\bm{y},\bm{\theta},\sigma_{\mathrm{N}}^{2}\sim\mathcal{N}(\bm{\mu}_{\mathrm{C}},\mathbf{Q}_{\mathrm{C}}^{-1}),$ (14)
@@
$\mathbf{Q}_{\mathrm{C}}=\mathbf{Q}_{z}+\sigma^{-2}_{\mathrm{N}}\mathbf{S}^{\top}\mathbf{S},$ (15)
@@
$\bm{\mu}_{\mathrm{C}}=\sigma^{-2}_{\mathrm{N}}\mathbf{Q}_{\mathrm{C}}^{-1}\mathbf{S}^{\top}\bm{y}.$ (16)
Using this conditional distribution, the point prediction of the true value $y^{*}$ at the unobserved location $\bm{s}^{*}\in\mathcal{D}$ is given by
$\mathbb{E}\left[\bm{y}^{*}|\bm{y},\sigma_{\mathrm{N}}^{2},\bm{\theta}\right]=\mathbf{S}_{\mathrm{P}}\bm{\mu}_{\mathrm{C}},$
where $\mathbf{S}_{\mathrm{P}}$ is the corresponding design matrix for the prediction location. The predictive distribution $\bm{y}^{*}|\bm{y},\sigma_{\mathrm{N}}^{2},\bm{\theta}$ is Gaussian, and the variance can be found either through algorithms that compute a partial inverse of $\mathbf{Q}_{\mathrm{C}}$ at a similar computational complexity as a Cholesky factorization (Rue and Held,, 2010), estimated through simulation, or through matrix free algorithms such as in Clarotto et al., (2024). When forecasts are desired, one can either include additional unobserved time steps in the discretization, or simulate as above until the last time point with data and then simulate forward in time using the forward dynamics in Equation (12).
@@ 3.5 Parameter inference
$\log\pi(\bm{\theta},\sigma^{2}_{\mathrm{N}}|\bm{y})$ $=\mathrm{Const}+\log\pi(\bm{\theta},\sigma^{2}_{\mathrm{N}})+\frac{1}{2}\log|\mathbf{Q}_{z}|-\frac{1}{2}\log|\mathbf{Q}_{\mathrm{C}}|+$ (17)
$\phantom{=}-\frac{n}{2}\log\sigma^{2}_{\mathrm{N}}-\frac{1}{2}\bm{\mu}_{\mathrm{C}}^{\mathrm{T}}\mathbf{Q}_{z}\bm{\mu}_{\mathrm{C}}-\frac{1}{2\sigma^{2}_{\mathrm{N}}}(\bm{y}-\mathbf{S}\bm{\mu}_{\mathrm{C}})^{\top}(\bm{y}-\mathbf{S}\bm{\mu}_{\mathrm{C}}),$
where we use $\pi(\bm{\theta},\sigma_{\mathrm{N}}^{2})\propto 1$ in this paper since we are emulating densely observed numerical model output and do not need extra regularization. For the derivation of the log-likelihood, we refer the reader to Berild and Fuglstad, (2023, Appendix A.4).
Inference is done using an optimization algorithm on the log-posterior to find the parameters that maximize it. To aid the optimization, the gradient of the log-posterior is used. Calculating the analytic gradient of the log-posterior requires computing $\mathrm{Tr}((\mathbf{Q}_{z}^{-1}-\mathbf{Q}_{\mathrm{C}}^{-1})\frac{\partial\mathbf{Q}_{z}}{\partial\theta_{i}})$, where $\mathrm{Tr}(\cdot)$ denotes the trace of the matrix. Fuglstad et al., 2015b took advantage of sparsity by using an algorithm that computes a partial inverse only for the non-zero elements (Rue and Held,, 2010), but computations are still expensive. We use a stochastic trace estimator (STE), through the Hutchinson estimator (Hutchinson,, 1990), to approximate the trace by
$\widehat{\mathrm{Tr}(\mathbf{A})}=\widehat{\mathrm{E}\left[\bm{\xi}^{\top}\mathbf{A}\bm{\xi}\right]}=\frac{1}{n}\sum_{i=1}^{n}\bm{\xi}_{i}^{\top}\mathbf{A}\bm{\xi}_{i},$
where $n$ is an integer, and all elements of the vectors $\bm{\xi}_{1},\ldots,\bm{\xi}_{n}$ are independent samples from the Rademacher distribution, i.e., $\xi_{i,j}=\pm 1$ with equal probability for all $i$ and $j$. This approximation was also used in Pereira et al., (2022) and Clarotto et al., (2024), and is much faster than analytic calculations as it only requires $n$ matrix-vector products and solves.
The drawback of stochastic optimization is that the gradient is stochastic, and this can lead to erratic behavior in the optimization, but it is often used because the computational speedup is so significant that the convergence is still faster. There is a vast literature on stochastic gradient methods, and we refer the reader to Ruder, (2017) for an overview. Many of these methods tackle the problem of noisy gradients and use techniques such as momentum, adaptive learning rates, and averaging of gradients to reduce the noise. In this work, we use the Adam optimizer (Kingma and Ba,, 2017) which is a popular choice for optimization of neural networks and other large models.
@@ 4.1 True process and observation model
For simplicity, the dampening coefficient is constant $\kappa(\cdot)\equiv\exp(-2)$, the isotropic diffusion parameter is $\gamma(\cdot)\equiv\exp(-1)$, and the scaling parameter for the forcing is $\tau=\exp(-4)$. This defines the true spatio-temporal process $\{u_{\mathrm{true}}(\bm{s},t):\bm{s}\in\mathcal{D},t\in\mathcal{T}\}$, which we call NStat-True.
At the boundaries, we impose zero flow going in and out of the domain. As a consequence, diffusion inflates variance close to the boundaries, and variance is inflated near the boundary if there is strong advection into the boundary and variance is deflated if there is strong advection away from the boundary. To make behaviour more realistic within $\mathcal{D}$ and avoid large increases in variance, we add a buffer of $5$ grid cells to each side of the domain where we taper advection linearly to zero in the buffer. The new domain $\mathcal{D}_{\mathrm{B}}$ then has $M=60$ and $N=60$ grid cells.
@@
$y_{i}=u_{\mathrm{true}}(\bm{s}_{i},t_{i})+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma_{\mathrm{N}}^{2}),$ (20)
for all cells $i$. Here, the noise is i.i.d. Gaussian with variance $\sigma_{\mathrm{N}}^{2}=0.001$. Figure 3 shows one realization, marginal variance, and spatio-temporal correlation with the marked location.
@@ 4.2 Goals, and training and test data
We aim to mimic the emulation of a numerical model that has been densely observed with a high signal-to-noise ratio. We generate 20 spatio-temporal realizations of NStat-True that is used for training and testing. In each of the 20 realizations, the training data excludes the rectangular area $[6.1,11.6]\times[3.7,9.5]\subset\mathcal{D}$ (rounded to the first decimal), and the test data contains only the rectangular area. One realization of the training data is demonstrated in Figure 3a. The goal is to assess the ability to reconstruct the missing rectangle, and to forecast for different lags forward in time.
There are $n_{\mathrm{train}}$ spatial training locations and $n_{\mathrm{test}}$ spatial test locations, and we can denote the training data and test data as $\bm{y}^{\mathrm{train}}_{t,r}=(y^{\mathrm{train}}_{1,t,r},\ldots,y^{\mathrm{train}}_{n_{\mathrm{train}},t,r})^{\mathrm{T}}$ and $(y^{\mathrm{test}}_{1,t,r},\ldots,y^{\mathrm{test}}_{n_{\mathrm{test}},t,r})^{\mathrm{T}}$, respectively, for time points $t=1,\ldots,12$ and realizations $r=1,\ldots,20$. Parameters $\hat{\bm{\theta}}$ for each candidate model are estimated using all training data $\bm{y}^{\mathrm{train}}_{t,r}$, $t=1,\ldots 12$ and $r=1,\ldots,20$. We explain for a generic model and drop labels to distinguish between models for ease of presentation.
For each realization $r=1,\ldots,20$ for each time point $t=1,\ldots,12$, we consider the $k$-step ahead point predictions
$\hat{\bm{y}}^{\mathrm{test}}_{k,t,r}=\mathrm{E}[\bm{y}^{\mathrm{test}}_{t+k,r}|\bm{y}^{\mathrm{train}}_{t,r},\hat{\bm{\theta}}],\quad k=0,\ldots,12-t.$
Note that $k=0$ is a reconstruction of the missing rectangle for a given time point. The point predictions are evaluated using root mean square error (RMSE)
$\mathrm{RMSE}_{k,t,r}=\sqrt{\frac{1}{n_{\mathrm{test}}}||\bm{y}^{\mathrm{test}}_{t+k,r}-\hat{\bm{y}}^{\mathrm{test}}_{k,t,r}||^{2}},\quad t=1,\ldots,12-k,\quad r=1,\ldots 20,$
for $k=0,\ldots,12$, where |||||$\cdot$||| | ⋅ | | is the standard Euclidean distance. This gives 240 RMSE values for $k=0$ and 20 RMSE values for $k=12$.
We also assess the predictive distributions using the continuous ranked probability score (CRPS, Gneiting and Raftery, (2007)),
$\mathrm{CRPS}_{k,t,r}=\frac{1}{n_{\mathrm{test}}}\sum_{i=1}^{n_{\mathrm{test}}}\sigma_{i,k,t,r}(z_{i,k,t,r}(2\Phi(z_{i,k,t,r})-1)+2\phi(z_{i,k,t,r})-\pi^{-1/2}),$
where $\sigma_{i,k,t,r}$ is the prediction variance for test location $i$ based on the predictive distribution $\bm{y}^{\mathrm{test}}_{t+k,r}|\bm{y}^{\mathrm{train}}_{t,r},\hat{\bm{\theta}}$, and $z_{i,k,t,r}$ is the difference between element $i$ in $\bm{y}^{\mathrm{test}}_{t+k,r}$ and element $i$ of the point forecast $\hat{\bm{y}}^{\mathrm{test}}_{k,t,r}=\mathrm{E}[\bm{y}^{\mathrm{test}}_{t+k,r}|\bm{y}^{\mathrm{train}}_{t,r},\hat{\bm{\theta}}]$, divided by $\sigma_{i,k,t,r}$. Here $\Phi(\cdot)$ is the cumulative distribution function of the standard normal distribution, and $\phi(\cdot)$ is the density function of the standard normal distribution.
For each $k=0,\ldots,12$, we summarize RMSE and CRPS across all $t=1,\ldots,12-k$ and $r=1,\ldots,20$ using the average and the empirical standard deviation.